A certain recipe asks for 5/7 cup of granola. How much is needed to make 3/2 of a recipe.
Answer by stanbon(75887)

A certain recipe asks for 5/7 cup of granola.
How much is needed to make 3/2 of a recipe.
------
Use a proportion::
x/(3/2) = (5/7)/1
------
x = (3/2)(5/7)
------
x = 15/14
==============
Cheers,
Stan H.
==============